OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (anonymous):

First I would jump to part two since knowing the right answer will help you answer part 1. Writing it in number form rather than written form will help so we will start there: \[3\sqrt{2x}+x \sqrt{8x}-5\sqrt{18x}=? \] \[nakim's answer=-10\sqrt{2x}\] Now lets simplify the equation and find our own answer: Since \[\sqrt{2x}\] is in the simplest form lets try to get that on each term of the equation. \[3\sqrt{2x}+x \sqrt{4}\sqrt{2x}-5\sqrt{9}\sqrt{2x}=? \] Then you can simplify from there to find the answer. Let me know if you need more help.